OpenStudy (theresa.womack):

What are the complex solutions for x^2=4x-29?

OpenStudy (anonymous):

2 plus minus 5i

OpenStudy (theresa.womack):

Thank you!

OpenStudy (anonymous):

\[ x^2-4x+29=0\\ \text{Use the quadratic formula:}\\ \dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\\ \dfrac{4\pm\sqrt{16-4\cdot29}}{2}\\ \dfrac{4\pm\sqrt{-100}}{2}\\ \dfrac{4\pm10i}{2}\\ 2\pm5i \]